OpenStudy (anonymous):

if f(x) = sec x and g(x)= sin x , which statement is true, a) (f×g)(x)= cosx b) (f×g)(x)= cotx c) (f×g)(x)= cscx d) (f×g)(x)= tanx

OpenStudy (unklerhaukus):

\[(f×g)(x)=f(x)g(x)=\]

OpenStudy (unklerhaukus):

simplify using \[\sec=\frac1\cos\]

OpenStudy (anonymous):

OH WAIT.... got it ..

OpenStudy (anonymous):

sec x = 1/cosx 1/cosx (sinx) = sinx/cosx = tan x

terenzreignz (terenzreignz):

Nice ^_^

OpenStudy (unklerhaukus):

Excellent work @alexeis_nicole \[\boxed{\huge\color{red}\checkmark}\]

OpenStudy (anonymous):

thanks! sorry if i didn't get that quicker than i should have... haven't stopped studying since yesterday morning. i know i should put this down but im actually terrified to